# Video Anomaly Detection in 10 Years: A Survey and Outlook

### v1

@@ Abstract
Video Anomaly Detection, Weak Supervision, Reconstruction-based Techniques, Vision-Language Models, Video Surveillance.
@@ I Introduction
While the majority of deep learning-based VAD systems rely on supervised learning paradigms, recent years have witnessed a surge in interest in exploring novel approaches, including weakly supervised, self-supervised, and unsupervised methods [9, 10, 11]. These alternative approaches present promising solutions to the challenges encountered by conventional VAD methods, such as the requirement for extensively annotated datasets and the complexity of capturing intricate spatiotemporal patterns. Through the utilization of deep learning techniques, researchers aspire to cultivate robust and efficient VAD systems adept at navigating diverse real-world scenarios and applications [12, 1]. Similarly, several research challenges are either not thoroughly explored or remain unexplored. For instance, a crucial consideration is the quality and diversity of available datasets. The development of emerging comprehensive datasets, such as the UCF Crime dataset [11], XD-Violence dataset [13], and ShanghaiTech dataset [14], plays a pivotal role in this progress. These datasets encompass various types of anomalies, significantly contributing to the enhancement of VAD techniques.
@@
In VAD, videos consist of sequences of frames, resulting in complex, high-dimensional spatiotemporal data. Detecting anomalies within such complex data necessitates methods capable of effectively capturing spatial, temporal, spatiotemporal, and textual features. To address these challenges, numerous VAD methods and deep feature extractors have been introduced in academic research, which has played a significant role in advancing the current state-of-the-art. It is also worth mentioning the importance of choosing a correct loss function based on the type of VAD paradigm. Hence, loss functions and regularization techniques are another challenge in the VAD problem. Another significant challenge lies in the diverse paradigms including self-supervised, weakly supervised, fully supervised, and vision-language models utilized to address VAD problem. These paradigms can be categorized into classical and deep learning. The classical approaches employ handcrafted features including spatiotemporal gradient [15], Histograms of Oriented Gradients (HOG) features [16, 17], and Histograms of Optical Flows (HOF) [18] within a temporal cuboid. These features are selected for their effectiveness in capturing appearance and motion information in a spatiotemporal context. The deep learning approaches are more powerful in terms of rich feature representation and end-to-end learning and have gained much popularity in the past decade due to advancements in different learning models. Specifically, these approaches leverage powerful feature extractors to capture meaningful spatiotemporal features. Examples include convolutional neural networks [19], autoencoders [20], GANs [21], vision transformers [22, 23], and vision language models [24, 25].
@@
In addition to the aforementioned inherent challenges within the VAD problem, each paradigm comes up with diverse loss functions, spatiotemporal regularization/constraints, and anomaly score prediction components. We also identify these different modules in this survey and provide take-home messages for each of the abovementioned challenges. Figure  1 shows the performance variation of representative deep learning-based VAD methods on two publicly available benchmark datasets including UCF-Crime [11] and ShanghaiTech [14]. The performance is reported in terms of the area under the ROC curve (AUC%). The performance improvement trend in this figure underscores a significant advancement in deep learning methodologies over the past decade, with the most notable improvement achieved by utilizing the recently proposed vision-language model [26].
@@
Figure 1: Performance improvement from 2017 until 2023 on two popular benchmarks. Performance is measured by the area under the ROC curve (AUC%). Note that some of the models were developed before the datasets were created but were used after the creation by other researchers such as [15, 16]. The other proposed models are : [27, 28, 29, 30, 31, 32, 23, 33, 26]
@@
Motivation: Despite the growing interest in VAD and the proliferation of deep learning-based approaches, there remains a need for a comprehensive survey that explores the latest developments in this field. Existing surveys often focus on traditional VAD methods and may not adequately cover the emerging trends and methodologies such as vision-language models (VLMs) [34, 35, 33, 36]. By delving into VLMs and deep learning-based VAD, including supervised, weakly supervised, self-supervised, and unsupervised approaches, this survey aims to provide a thorough understanding of the state-of-the-art techniques and their potential applications. Our main focus lies on deep learning-based solutions for VAD problems. Particularly, we explore emerging VAD paradigms that employ different datasets, feature extractions, loss functions, and spatiotemporal regularization. In addition, we offer a comprehensive analysis of over 50 different methodologies employed in VAD, with a specific focus on the potential of textual features extracted with vision language models. In this survey, our main contributions are summarized as follows:
@@
The structure of this paper is as follows: Section II outlines the methodology used to select the research studies included in this review over the past decade. Section III examines previous surveys conducted in the field of video anomaly detection. Section IV presents the formulation of the video anomaly detection problem. Section V presents core VAD challenges. Section VI outlines experimentation and evaluation protocols used. Section VII provides a comparative analysis through quantitative and qualitative comparisons of state-of-the-art models. Section VIII provides visualization of bibliometric networks for thematic analysis. Finally, we conclude our work and present additional future directions in Section IX.
@@ III Previous Literature Review
In a different work, Nayak et al. [7] categorize learning frameworks into four main categories: supervised, unsupervised, semi-supervised, and active learning. In the context of deep-learning-based VAD, state-of-the-art methods fall into several distinct categories, including Trajectory-based methods, Global pattern-based methods, Grid pattern-based methods, Representation learning models, Discriminative models, Predictive models, Deep generative models, Deep one-class deep neural networks, and Deep hybrid models.
@@
Moreover, the research provides a comprehensive analysis of performance evaluation methodologies, covering aspects such as the choice of datasets, computational infrastructure, evaluation criteria, and performance metrics. Another recent work by Pang et al. [4] focused on deep learning techniques for anomaly detection and explored various challenges within the anomaly detection (AD) problem. These challenges included issues such as class imbalance in the data, complex anomaly detection, the presence of noisy instances in weakly supervised AD methods, and more. They discussed how deep learning methods offer solutions to these diverse challenges. To structure their analysis, Pang et al. introduced a hierarchical taxonomy for categorizing deep anomaly detection methods. This taxonomy comprised three principal categories: deep learning for feature extraction, learning feature representations of normality, and end-to-end anomaly score learning. Additionally, they provided 11 fine-grained subcategories from a modeling perspective.
@@
Previous research efforts have overlooked the critical importance of employing diverse feature extractors within deep learning models, especially in the context of video anomaly detection. This oversight becomes glaringly apparent when considering emerging trends such as the widespread adoption of transformers and vision-language pretraining models within video anomaly detection. These innovations have significantly influenced the overall performance of deep learning models. In our survey paper, we aim to address these research gaps that have not been adequately outlined in previous surveys.
Furthermore, our study strongly emphasizes the latest trends in VAD, ensuring that our analysis remains current and relevant in the rapidly evolving landscape of anomaly detection research. We go beyond simply summarizing existing literature; instead, we explore recent advancements that build upon the foundation laid by earlier survey papers in this field. By synthesizing this wealth of information, we provide readers with a comprehensive overview of the state-of-the-art in VAD.
@@ IV Defining the Video Anomaly Detection Problem
Anomalies within video data encompass events or behaviors that notably diverge from anticipated or regular patterns. The primary goal of video anomaly detection is to devise and implement resilient algorithms and models capable of autonomously identifying and flagging these anomalies in real time. This involves converting raw video data into interpretable feature representations utilizing robust feature extractors adept at capturing both spatial and temporal characteristics. Additionally, it necessitates the selection of appropriate algorithms or techniques and the establishment of effective evaluation metrics to assess detection performance accurately.
@@ V Analyzing Video Anomaly Detection: A Systematic Approach
Our exploration begins with the utilization of diverse datasets and a spectrum of feature extraction techniques, encompassing CNN, 3D CNN, Autoencoders, GANs, LSTMs, Transformers, vision language features, and Hybrid feature extractors. These methodologies are applied to extract spatial, temporal, spatiotemporal, and textual features. Additionally, the section will address various learning approaches, including supervised, self-supervised, and reconstruction-based strategies, which often fall within the realm of unsupervised or one-class classification methods, alongside weakly supervised and predictive models. We will also examine into aspects such as loss functions, regularization methods, and anomaly score computation. Our objective is to provide a comprehensive analysis of these challenges and review the solutions proposed by researchers in this field.
Furthermore, our exploration will encompass various learning and supervision strategies, including supervised methods, self-supervised techniques, and unsupervised techniques (often classified as reconstruction-based or one-class classification approaches), alongside weakly supervised and prediction methods. Additionally, we will illuminate the significance of loss functions, regularization techniques, and anomaly score computation. Evaluation protocols for models are discussed in section VI-B. In the upcoming sections, we will explore these challenges further, examining the ways they have been approached.
@@ V-A Datasets Building and Selection
The field of Video Anomaly Detection (VA relies significantly on publicly available datasets which are used for testing and benchmarking the proposed models. In this section, we present an overview of the common datasets utilized in the field of VAD, each carefully curated to facilitate the study of anomalies in various scenarios. These datasets encompass a wide range of scenes, offering diverse challenges for anomaly detection. We will investigate the number of videos contained in each dataset, the specific types of scenes they cover, and the anomalies present.
@@ V-A2 UCSD Pedestrian
V-A3 UCF-Crime dataset
@@ V-A4 CUHK Avenue
V-A5 ShanghaiTech
The ShanghaiTech Campus dataset [14] stands as a substantial contribution to the realm of anomaly detection, presenting a vast and diverse collection of data. Across its 13 scenes, characterized by intricate light conditions and varied camera angles, the dataset encapsulates a comprehensive array of challenging scenarios. Notably, it boasts an extensive compilation of over 270,000 training frames, alongside 130 abnormal events meticulously annotated at the pixel level, facilitating precise evaluation and analysis.
@@ V-A6 XD-Violence
V-A7 NWPU Campus dataset
@@ V-A8 Discussion on Datasets
  • A recurring theme in multiple datasets is the limited number of anomalous events contained in the dataset. Some datasets contained as little as twelve anomalous event types (UCSD Ped2 [39]) in comparison to other datasets that featured more than 130 types of anomalies (ShanghaiTech [14]). This constraint lessens the models’ efficacy in practical applications by limiting their capacity to learn a broad range of potential abnormal behaviors.

### v2

@@ Abstract
Reconstruction-based Techniques, Video Anomaly Detection, Vision-Language Models, Video Surveillance, Weak Supervision.
[…]
=-15pt
@@ I Introduction
Alongside the deep learning-based VAD systems that rely on supervised learning paradigms, recent years have witnessed a surge in exploring novel approaches, including weakly supervised, self-supervised, and unsupervised methods [9, 10, 11]. These alternative approaches present promising solutions to the challenges encountered by conventional VAD methods, such as the requirement for extensively annotated datasets and the complexity of capturing intricate spatiotemporal patterns. Through the utilization of deep learning techniques, researchers aspire to cultivate robust and efficient VAD systems adept at navigating diverse real-world scenarios and applications [12, 1]. Similarly, several research challenges are either not thoroughly explored or remain unexplored. For instance, a crucial consideration is the quality and diversity of available datasets. The development of emerging comprehensive datasets, such as the UCF Crime dataset [11], XD-Violence dataset [13], and ShanghaiTech dataset [14], plays a pivotal role in this progress. These datasets encompass various types of anomalies, significantly contributing to the enhancement of VAD techniques.
@@
In VAD, videos consist of sequences of frames, resulting in complex, high-dimensional spatiotemporal data. Detecting anomalies within such complex data necessitates methods capable of effectively capturing spatial, temporal, spatiotemporal, and textual features. To address these challenges, numerous VAD methods and deep feature extractors have been introduced in academic research, which has played a significant role in advancing the current state-of-the-art. It is also worth mentioning the importance of choosing a correct loss function based on the type of VAD paradigm. Hence, loss functions and regularization techniques are another challenge in the VAD problem. Another significant challenge lies in the diverse paradigms including self-supervised, weakly supervised, fully supervised, and unsupervised models utilized to address the VAD problem.
These paradigms can be categorized into classical and deep learning. The classical approaches employ handcrafted features including spatiotemporal gradient [15], Histograms of Oriented Gradients (HOG) features [16, 17], and Histograms of Optical Flows (HOF) [18] within a temporal cuboid. These features are selected for their effectiveness in capturing appearance and motion information in a spatiotemporal context.
The deep learning approaches are more powerful in terms of rich feature representation and end-to-end learning and have gained much popularity in the past decade due to advancements in different learning models. Specifically, these approaches leverage powerful feature extractors to capture meaningful spatiotemporal features. Examples include convolutional neural networks [19], autoencoders [20], GANs [21], vision transformers [22, 23], and vision language models [24, 25].
@@
In addition to the aforementioned inherent challenges within the VAD problem, each paradigm comes up with diverse loss functions, spatiotemporal regularization/constraints, and anomaly score prediction components. We also identify these different modules in this survey and provide take-home messages for each of the abovementioned challenges. Figure  LABEL:fig:introchart shows the performance variation of representative deep learning-based VAD methods on two publicly available benchmark datasets including UCF-Crime [11] and ShanghaiTech [14]. The performance is reported in terms of the area under the ROC curve (AUC%). The performance improvement trend in this figure underscores a significant advancement in deep learning methodologies over the past decade, with the most notable improvement achieved by utilizing the recently proposed vision-language-based model [26].
[…]
@@
[t!](topskip=0pt, botskip=0pt, midskip=0pt)[width=0.999]Figures/introchart.pngPerformance improvement from 2017 until 2023 on two popular benchmarks. Performance is measured by the area under the ROC curve (AUC%). Note that some of the models were developed before the datasets were created but were used after the creation by other researchers such as [15, 16], the other proposed models are: [27],[28],[29],[30],[31],[32],[23],[33],[26].
@@
Motivation: Despite the growing interest in VAD and the proliferation of deep learning-based approaches, there remains a need for a comprehensive survey that explores the latest developments in this field. Existing surveys often focus on traditional VAD methods and may not adequately cover the emerging trends and methodologies such as vision-language models (VLMs) [34, 35, 33, 36, 26]. By delving into VLMs and deep learning-based VAD, including supervised, weakly supervised, self-supervised, and unsupervised approaches, this survey aims to provide a thorough understanding of the state-of-the-art techniques and their potential applications. Our main focus lies on deep learning-based solutions for VAD problems. Particularly, we explore emerging VAD paradigms that employ different datasets, feature extractions, loss functions, and spatiotemporal regularization. In addition, we offer a comprehensive analysis of over 50 different methodologies employed in VAD, with a specific focus on the potential of textual features extracted with vision language models. In this survey, our main contributions are summarized as follows:
@@
The structure of this paper is as follows: Section II outlines the methodology used to select the research studies included in this review over the past decade. Section III examines previous surveys conducted in the field of video anomaly detection. Section IV presents the formulation of the video anomaly detection problem. Section V presents a systematic approach and taxonomy for analyzing the VAD problem, and then the core challenges including the datasets V-A, feature extraction V-B, supervision schemes V-C, loss functions V-D, regularization techniques V-E and the anomaly score V-F. Section VI outlines datasets guidelines and evaluation protocols used. Section VII provides a comparative analysis through quantitative and qualitative comparisons of state-of-the-art models. Section VIII provides visualization of bibliometric networks for thematic analysis. Finally, we conclude our work and present additional future directions in Section IX
@@ III Related work
In a different work, Nayak et al. [7] categorize learning frameworks into four main categories: supervised, unsupervised, semi-supervised, and active learning. In the context of deep-learning-based VAD, state-of-the-art methods fall into several distinct categories, including Trajectory-based methods, Global pattern-based methods, Grid pattern-based methods, Representation learning models, Discriminative models, Predictive models, Deep generative models, Deep one-class deep neural networks, and Deep hybrid models. Moreover, the research provides a comprehensive analysis of performance evaluation methodologies, covering aspects such as the choice of datasets, computational infrastructure, evaluation criteria, and performance metrics.
@@
Another work by Pang et al. [4] focused on deep learning techniques for anomaly detection and explored various challenges within the anomaly detection (AD) problem. These challenges included issues such as class imbalance in the data, complex anomaly detection, the presence of noisy instances in weakly supervised AD methods, and more. They discussed how deep learning methods offer solutions to these diverse challenges. To structure their analysis, Pang et al. introduced a hierarchical taxonomy for categorizing deep anomaly detection methods. This taxonomy comprised three principal categories: deep learning for feature extraction, learning feature representations of normality, and end-to-end anomaly score learning. Additionally, they provided 11 fine-grained subcategories from a modeling perspective.
In their research, Mohammad Baradaran and Robert Bergevin [38] delve deeply into the semi-supervised VAD approaches, focusing on scenarios where labeled anomaly data is limited. They emphasize the role of feature extractors in these contexts, highlighting their ability to differentiate intricate patterns within video data by capturing crucial spatial and temporal details. These feature extractors are pivotal for detecting anomalies in semi-supervised tasks, where the model learns predominantly from an extensive set of normal data. The authors conduct an experimental analysis to shed light on the strengths and weaknesses of various VAD methods. They categorize the DL semi-supervised approaches into six distinct types: reconstruction, prediction, memorization, object-centric, segmentation, and multi-task learning-based methods. For each category, they provide a comprehensive examination of the strengths and shortcomings, particularly focusing on the effectiveness of different feature extraction techniques.
Another recently published survey paper by Nomica et al. [39] provides an in-depth analysis of machine learning techniques for detecting anomalies in video surveillance systems. It categorizes these methods into supervised, semi-supervised, and unsupervised approaches, highlighting their strengths, weaknesses, and applicability. However, it does not address the critical differences between feature types—such as temporal, spatial, textual, and hybrid features. These differences significantly influence the choice of feature extractors, impacting the effectiveness of the detection models. Additionally, it significantly overlooks the topic of Vision-Language Models as feature extractors. In contrast, our work provides an in-depth analysis that encompasses all these aspects, offering a comprehensive understanding of their impact on anomaly detection models.
A very recently published survey paper in 2024 by Yang et al. [40] categorizes the VAD approaches into unsupervised, Weakly-supervised, fully unsupervised, and supervised VAD. It highlights their strengths, such as improved feature extraction and detailed object analysis, while also noting the importance of handling spatiotemporal features and illumination changes. Although the survey provides an extensive comparison of different VAD methodologies, it notably omits the discussion of Vision-Language Models as feature extractors, a growing area of interest in the field.
@@
Previous research efforts have overlooked the critical importance of employing diverse feature extractors within deep learning models. This oversight becomes apparent when considering emerging trends such as the widespread adoption of transformers and vision-language pretraining models within video anomaly detection. These innovations have significantly influenced the overall performance of deep learning models. In our survey paper, we aim to address these research gaps that have not been adequately outlined in previous surveys.
@@
III-A Contributions
In this survey, our main contributions compared to existing survey papers are summarized as follows:
  • This work defines a clear and comprehensive problem formulation of the VAD problem specifically in the context of supervised learning, where frame-level labels are available.
  • To the best of our knowledge, this is the first survey to highlight the emerging importance of integrating Vision-Language Models (VLMs) as feature extraction in VAD. We explore how VLMs can significantly enhance model performance by effectively combining visual and textual data to better understand and detect anomalies.
  • The paper is organized to serve as a detailed guide for readers and researchers new to the field of VAD. We provide foundational knowledge and a structured approach to navigate the complexities of VAD research as shown in Figure 1.
  • This work presents a well-structured discussion starting from the selection of datasets and the types of features that are critical for VAD. This includes a focus on textual features and deep feature extractors. We also explore the various learning and supervision paradigms, including supervised, self-supervised, weakly supervised, and unsupervised/reconstruction approaches, detailing their respective advantages and disadvantages.
  • This work provides a taxonomy of video anomaly detection that is systematically categorized into two main dimensions: learning and supervision schemes, and feature extraction as shown in Figure 2
  • This work provides insights into selecting appropriate loss functions and regularization techniques, which are crucial for optimizing the performance of VAD models.
  • A comprehensive guideline is provided for choosing the most suitable datasets and evaluation metrics for experimental purposes, ensuring that researchers can effectively assess and compare their VAD methods.
  • This work offers a detailed comparative analysis of the state-of-the-art (SOTA) VAD methods, both quantitatively and qualitatively. This analysis helps in understanding the current landscape and performance benchmarks in the field.
  • Finally, this work provides an extensive discussion on potential future research directions in VAD. This includes exploring new technologies, methodologies, and application areas, and guiding researchers toward promising avenues for further investigation.
@@ IV Defining the Video Anomaly Detection Problem
Anomalies within video data encompass events or behaviors that notably diverge from anticipated or regular patterns. The primary goal of video anomaly detection is to devise and implement resilient algorithms and models capable of autonomously identifying and flagging these anomalies in real-time. This involves converting raw video data into interpretable feature representations utilizing robust feature extractors adept at capturing both spatial and temporal characteristics. Additionally, it necessitates the selection of appropriate algorithms or techniques and the establishment of effective evaluation metrics to assess detection performance accurately.
@@ V Analyzing Video Anomaly Detection: A Systematic Approach
Our exploration begins with the utilization of diverse datasets and a spectrum of feature extraction techniques. These methodologies are applied to extract spatial, temporal, spatiotemporal, and textual features. Furthermore, our exploration will encompass various learning and supervision strategies, including supervised methods, self-supervised techniques, and unsupervised techniques (often classified as reconstruction-based or one-class classification approaches), alongside weakly supervised and prediction methods. Additionally, we will illuminate the significance of loss functions, regularization techniques, and anomaly score computation. Evaluation protocols for models are discussed in section VI-B. In the upcoming sections, we will explore these challenges further, examining the ways they have been approached.
Taxonomy of Video Anomaly Detection
The taxonomy of video anomaly detection, as shown in Figure 2, is systematically categorized into two main dimensions: Learning and Supervision Schemes, and Feature Extraction. The Learning and Supervision Schemes dimension includes Supervised, Self-Supervised, Weakly Supervised (such as Multiple Instance Learning), and Unsupervised methods. Unsupervised methods are further divided into One-Class Classification, Reconstruction, and Future Frame Prediction approaches. The Feature Extraction dimension involves Deep Feature Extractors, including CNNs, Autoencoders, GANs, Sequential Deep Learning models (like LSTMs and Vision Transformers), Vision Language models, and Hybrid models. Additionally, it covers different types of features such as Spatial, Temporal, SpatioTemporal, and Textual.
Figure 2: Taxonomy of Video Anomaly Detection.
@@ V-A Datasets Building and Selection
The field of Video Anomaly Detection relies significantly on publicly available datasets which are used for testing and benchmarking the proposed models. In this section, we present an overview of the common datasets utilized in the field of VAD, each carefully curated to facilitate the study of anomalies in various scenarios. These datasets encompass a wide range of scenes, offering diverse challenges for anomaly detection. We will investigate the number of videos contained in each dataset, the specific types of scenes they cover, and the anomalies present.
@@ V-A2 UCSD Pedestrian
V-A3 Street Scene dataset
The Street Scene dataset proposed by Ramachandra et al. [43] is designed for video anomaly detection and comprises 46 training and 35 testing high-resolution (1280×720) video sequences. These sequences were captured using a USB camera positioned to overlook a two-lane street with bike lanes and pedestrian sidewalks during the daytime. This dataset presents a challenging environment due to the diverse range of activities captured, including cars driving, turning, stopping, and parking; pedestrians walking, jogging, and pushing strollers; and bikers riding in bike lanes. Additionally, the videos feature changing shadows, moving backgrounds such as flags and trees blowing in the wind, and occlusions from trees and large vehicles.
The dataset includes 56,847 frames for training and 146,410 frames for testing, extracted at a rate of 15 frames per second. It contains a total of 205 naturally occurring anomalous events. These anomalies range from illegal activities like jaywalking and illegal U-turns to uncommon events not present in the training set, such as pets being walked and a meter maid issuing tickets. This variety makes the Street Scene dataset a comprehensive and demanding resource for advancing video anomaly detection research. However, a notable challenge of this dataset is that it is a single-scene dataset where only the street scene is included. This prevents models trained only on this dataset from generalizing their anomaly detection capabilities to other scenes. On the other hand, this dataset could be beneficial for models that are specifically designed to work on single scenes such as EVAL [44]. The dataset is publicly available on: http://www.merl.com/demos/video-anomaly-detection
V-A4 UCF-Crime dataset
@@ V-A5 CUHK Avenue
V-A6 ShanghaiTech
The ShanghaiTech Campus dataset [14] stands as a substantial contribution to the field of anomaly detection, presenting a vast and diverse collection of data. Across its 13 scenes, characterized by intricate light conditions and varied camera angles, the dataset encapsulates a comprehensive array of challenging scenarios. Notably, it boasts an extensive compilation of over 270,000 training frames, alongside 130 abnormal event occurrences meticulously annotated at the pixel level, facilitating precise evaluation and analysis.
@@ V-A7 XD-Violence
V-A8 NWPU Campus dataset
@@
Street Scene [43] 2020 No 81 226 mins
jaywalking,
illegal U-turns, pets,
metermaid ticketing a car
2,509 57:43
Single geographic location,
single scene,
only outdoor,
@@ V-A9 Discussion on Datasets
  • A recurring theme in multiple datasets is the limited number of anomalous events contained in the dataset. Some datasets contained as little as three anomalous event types (strange action, wrong direction, abnormal object, CUHK Avenue [15]) in comparison to other datasets that featured up to 11 types of anomalies (ShanghaiTech [14]). This constraint lessens the models’ efficacy in practical applications by limiting their capacity to learn a broad range of potential abnormal behaviors.